OpenStudy (anonymous):

-2x+10=4 fill in the table x o 1 2 3 4 -2x10 0 is 10 what is 1,2,3,4

OpenStudy (anonymous):

you could change the equation to y=-2x+6. Plug in your x's to get the y-values

OpenStudy (anonymous):

so for 1: we get -2(1)+6= 4

OpenStudy (anonymous):

what next

OpenStudy (anonymous):

then just plug in the next numbers

OpenStudy (anonymous):

so you have 2, plug in 2 for x

OpenStudy (anonymous):

6

OpenStudy (anonymous):

you got 6?

OpenStudy (anonymous):

it's: -2(2)+6 = 2

OpenStudy (anonymous):

what would 3 and 4 be

OpenStudy (anonymous):

-2(3)+6 = 0; -2(4)+6 = -2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome